john invested $20 000 on 1/06/2009 at 9% p.a. interest compounded monthly.
a) what is the value of the investment by 1/06/2019
b) If he only takes the interest earned on this investment on this date and invests this at 10.5% p.a. (compounded continuously), how much can he expect back by 1/06/2025

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

it will be 20000(1+.09/12)^1120=20000*(2.45 but round at end)=$49027.14 by 1/6/19
the interest is $29027.14 and that is multiplied by e^(.105*6)or e^0.735 =$60535.58 by 1/6/25
That formula is P=Poe^rt
Rule of 72 would say the first would double in 8 years and be about 2.5 times
The second would double in under 7 years, which it has.

Answer by MathTherapy(10552)


john invested $20 000 on 1/06/2009 at 9% p.a. interest compounded monthly.
a) what is the value of the investment by 1/06/2019
b) If he only takes the interest earned on this investment on this date and invests this at 10.5% p.a. (compounded continuously), how much can he expect back by 1/06/2025
Correct answers: